# GaussianImage: 1000 FPS Image Representation and Compression by 2D Gaussian Splatting

### v3

@@ Abstract
Implicit neural representations (INRs) recently achieved great success in image representation and compression, offering high visual quality and fast rendering speeds with 10-1000 FPS, assuming sufficient GPU resources are available. However, this requirement often hinders their use on low-end devices with limited memory. In response, we propose a groundbreaking paradigm of image representation and compression by 2D Gaussian Splatting, named GaussianImage. We first introduce 2D Gaussian to represent the image, where each Gaussian has 8 parameters including position, covariance and color. Subsequently, we unveil a novel rendering algorithm based on accumulated summation. Remarkably, our method with a minimum of 3$\times$ lower GPU memory usage and 5$\times$ faster fitting time not only rivals INRs (e.g., WIRE, I-NGP) in representation performance, but also delivers a faster rendering speed of 1500-2000 FPS regardless of parameter size. Furthermore, we integrate existing vector quantization technique to build an image codec. Experimental results demonstrate that our codec attains rate-distortion performance comparable to compression-based INRs such as COIN and COIN++, while facilitating decoding speeds of approximately 1000 FPS. Additionally, preliminary proof of concept shows that our codec surpasses COIN and COIN++ in performance when using partial bits-back coding. Code will be available at https://github.com/Xinjie-Q/GaussianImage.
@@
footnotetext: ${}^{\ast}$qual Contribution.footnotetext: ${}^{\dagger}$his work was done when Xinjie Zhang and Xingtong Ge interned at SenseTime Research.footnotetext: ${}^{\dotplus}$orresponding Authors.
@@ 1 Introduction
To address these issues, we propose a new paradigm of image representation and compression, namely GaussianImage, using 2D Gaussian Splatting. Firstly, we adopt 2D Gaussians in lieu of 3D for a compact and expressive representation. Each 2D Gaussian is defined by 4 attributes (9 parameters in total): position, anisotropic covariance, color coefficients, and opacity. This modification results in a 6.5$\times$ compression over 3D Gaussians with equivalent Gaussian points, significantly mitigating storage demands of Gaussian representation. Subsequently, we advocate a unique rasterization algorithm that replaces depth-based Gaussian sorting and $\alpha$-blending with a accumulated summation process. This novel approach directly computes each pixel’s color from the weighted sum of 2D Gaussians, which not only fully utilizes the information of all Gaussian points covering the current pixel to improve fitting performance, but also avoids the tedious calculation of accumulated transparency to accelerate training and inference speed. More important, this summation mechanism allows us to merge color coefficients and opacity into a singular set of weighted color coefficients, reducing parameter count to 8 and further improving the compression ratio to 7.375$\times$. Finally, we transfer our 2D Gaussian representation into a practical image codec. Framing image compression as a Gaussian attribute compression task, we employ a two-step compression strategy: attribute quantization-aware fine-tuning and encoding. By applying 16-bit float quantization, 6-bit integer quantization [11], and residual vector quantization (RVQ) [74] to positions, covariance parameters, and weighted color coefficients, respectively, followed by entropy coding to eliminate statistical correlations among quantized values, we successfully develop the first image codec based on 2D Gaussian Splatting. As a preliminary proof of concept, the partial bits-back coding [61, 53] is optionally used to further improve the compression performance of our codec. Overall, our contributions are threefold:
@@
    We present a new paradigm of image representation and compression by 2D Gaussian Splatting. With compact 2D Gaussian representation and a novel rasterization method, our approach achieves high representation performance with short training duration, minimal GPU memory overhead and remarkably, 2000 FPS rendering speed.
@@
    Experimental results show that when compared with existing INR methods, our approach achieves a remarkable training and inference acceleration with less GPU memory usage while maintaining similar visual quality. When used as an efficient image codec, our approach achieves compression performance competitive with COIN [24] and COIN++ [23]. Comprehensive ablations and analyses demonstrate the effectiveness of each proposed component.
@@ 3.2 Accumulated Blending-based Rasterization
Figure 3: Compression pipeline of our proposed GaussianImage. After overfitting image, we can apply a two-step compression procedure to build an ultra-fast image codec.
@@ 3.3 Compression Pipeline
After overfitting the image, we propose a compression pipeline for image compression with GaussianImage. As shown in Fig. 3, our standard compression pipeline is composed of three steps: image overfitting, attribute quantization-aware fine-tuning, and attribute encoding. To achieve the best compression performance, partial bits-back coding [61, 53] is an optional strategy. Herein, we elucidate the compression process using our GaussianImage based on Cholesky factorization as an example.
@@
Entropy Encoding. After quantization-aware fine-tuning, we employ entropy coding to further reduce the size of the quantized Gaussian parameters. By exploiting the statistical correlation of character frequencies, we adopt asymmetric numeral systems (ANS) [21] to encode the quantized covariance parameters and the color codebook indices, respectively.
@@
[…]
Despite its theoretical efficacy, bits-back coding may not align with the objective of developing an ultra-fast codec due to its slow processing latency [38]. Consequently, we leave this part as a preliminary proof of concept on the best rate-distortion performance our codec can achieve, instead of a final result of our codec can achieve with 1000 FPS.
@@ 4.2 Image Representation
Figure 5: Subjective comparison of our method against JPEG and COIN on Kodak.
Table 2: Computational complexity of traditional and learning-based image codecs on DIV2K Dataset.
@@
JPEG [62] 0.3197 25.2920 0.9020 608.61 614.68
JPEG2000 [56] 0.2394 27.2792 0.9305 3.46 4.32
Ballé17 [5] 0.2271 27.7168 0.9508 21.23 18.83
Ballé18 [6] 0.2533 28.7548 0.9584 16.53 15.87
COIN [24] 0.3419 25.8012 0.8905 8.68e-6 166.31
Ours 0.3164 25.6631 0.9154 4.11e-3 942.50
@@ 4.3 Image Compression
Computational complexity. Table 2 reports the computational complexity of several image codecs on the DIV2K dataset, with learning-based codecs operating on an NVIDIA V100 GPU and traditional codecs running on an Intel Core(TM) i9-10920X processor at a base frequency of 3.50GHz in single-thread mode. Impressively, the decoding speed of our codec reaches 942 FPS, outpacing traditional codecs like JPEG, while also providing enhanced compression performance at lower bitrates. This establishes a significant advancement in the field of neural image codecs.
@@ 4.4 Ablation Study
Quantization strategies. Table 4 investigates the effect of different quantization schemes on image compression. Without the commitment loss $\mathcal{L}_{c}$ (V1), the absence of supervision for the RVQ codebook leads to significant deviations of the codebook vector from the original vector, adversely affecting reconstruction quality. Moreover, eliminating RVQ in favor of 6-bit integer quantization for color parameters (V2) resulted in a 6.5% increase in bitrate consumption when compared with our default solution. This outcome suggests that the color vectors across different Gaussians share similarities, making them more suitable for RVQ. Further exploration into the use of higher bit quantization (V3) reveals a deterioration in compression efficiency.
@@
(V1) w/o $\mathcal{L}_{c}$+w/ RVQ + 6bit -3.061 294.82 -0.0820 304.21
(V2) w/o $\mathcal{L}_{c}$+w/o RVQ + 6bit -0.176 6.50 -0.0035 8.49
(V3) w/o $\mathcal{L}_{c}$+w/o RVQ + 8bit -0.248 10.16 -0.0081 16.80
@@ 0.C.1 Implementation Details
Baselines. For SIREN [55] and WIRE [54], we implement them by using the open-source projecthttps://github.com/vishwa91/wire from WIRE. For I-NGP [49] and NeuRBF [18], we adopt the projecthttps://github.com/oppo-us-research/NeuRBF from NeuRBF. As for compression baselines, the implementation of COIN [24] uses their libraryhttps://github.com/EmilienDupont/coin. We evaluate the VAE-based codecs (Ballé17 [5], Ballé18 [6]) using the MSE-optimized models provided by CompressAI [9]. It is worth noting that during the inference of the INR methods, we sample all image coordinates at once to output the corresponding RGB values. This is the maximum inference speed that the INR methods can achieve when the GPU memory resources are sufficient.

### v4

@@ Abstract
Implicit neural representations (INRs) recently achieved great success in image representation and compression, offering high visual quality and fast rendering speeds with 10-1000 FPS, assuming sufficient GPU resources are available. However, this requirement often hinders their use on low-end devices with limited memory. In response, we propose a groundbreaking paradigm of image representation and compression by 2D Gaussian Splatting, named GaussianImage. We first introduce 2D Gaussian to represent the image, where each Gaussian has 8 parameters including position, covariance and color. Subsequently, we unveil a novel rendering algorithm based on accumulated summation. Remarkably, our method with a minimum of 3$\times$ lower GPU memory usage and 5$\times$ faster fitting time not only rivals INRs (e.g., WIRE, I-NGP) in representation performance, but also delivers a faster rendering speed of 1500-2000 FPS regardless of parameter size. Furthermore, we integrate existing vector quantization technique to build an image codec. Experimental results demonstrate that our codec attains rate-distortion performance comparable to compression-based INRs such as COIN and COIN++, while facilitating decoding speeds of approximately 2000 FPS. Additionally, preliminary proof of concept shows that our codec surpasses COIN and COIN++ in performance when using partial bits-back coding. Code is available at https://github.com/Xinjie-Q/GaussianImage.
@@ 1 Introduction
To address these issues, we propose a new paradigm of image representation and compression, namely GaussianImage, using 2D Gaussian Splatting. Firstly, we adopt 2D Gaussians in lieu of 3D for a compact and expressive representation. Each 2D Gaussian is defined by 4 attributes (9 parameters in total): position, anisotropic covariance, color coefficients, and opacity. This modification results in a 6.5$\times$ compression over 3D Gaussians with equivalent Gaussian points, significantly mitigating storage demands of Gaussian representation. Subsequently, we advocate a unique rasterization algorithm that replaces depth-based Gaussian sorting and $\alpha$-blending with a accumulated summation process. This novel approach directly computes each pixel’s color from the weighted sum of 2D Gaussians, which not only fully utilizes the information of all Gaussian points covering the current pixel to improve fitting performance, but also avoids the tedious calculation of accumulated transparency to accelerate training and inference speed. More important, this summation mechanism allows us to merge color coefficients and opacity into a singular set of weighted color coefficients, reducing parameter count to 8 and further improving the compression ratio to 7.375$\times$. Finally, we transfer our 2D Gaussian representation into a practical image codec. Framing image compression as a Gaussian attribute compression task, we employ a two-step compression strategy: attribute quantization-aware fine-tuning and encoding. By applying 16-bit float quantization, 6-bit integer quantization [11], and residual vector quantization (RVQ) [73] to positions, covariance parameters, and weighted color coefficients, respectively, we successfully develop the first image codec based on 2D Gaussian Splatting. As a preliminary proof of concept, the partial bits-back coding [60, 52] is optionally used to further improve the compression performance of our codec. Overall, our contributions are threefold:
@@
    We present a pioneering paradigm of image representation and compression by 2D Gaussian Splatting. With compact 2D Gaussian representation and a novel accumulated blending-based rasterization method, our approach achieves high representation performance with short training duration, minimal GPU memory overhead and remarkably, 2000 FPS rendering speed.
@@
    Experimental results show that when compared with existing INR methods, our approach achieves a remarkable training and inference acceleration with less GPU memory usage while maintaining similar visual quality. When used as an efficient image codec, our approach offers competitive compression performance comparable to COIN [23] and COIN++ [22]. Comprehensive ablations and analyses demonstrate the effectiveness of each proposed component.
@@ 3.2 Accumulated Blending-based Rasterization
Figure 3: Compression pipeline of our proposed GaussianImage. After overfitting image, we apply attribute quantization-aware fine-tuning to build an ultra-fast image codec. Partial bits-back coding is used to achieve the best compression performance.
@@ 3.3 Compression Pipeline
After overfitting the image, we propose a compression pipeline for image compression with GaussianImage. As shown in Fig. 3, our standard compression pipeline is composed of two steps: image overfitting and attribute quantization-aware fine-tuning. To achieve the best compression performance, partial bits-back coding [60, 52] is an optional strategy. Herein, we elucidate the compression process using our GaussianImage based on Cholesky factorization as an example.
@@
Despite its theoretical efficacy, bits-back coding may not align with the objective of developing an ultra-fast codec due to its slow processing latency [37]. Consequently, we leave this part as a preliminary proof of concept on the best rate-distortion performance our codec can achieve, instead of a final result of our codec can achieve with 2000 FPS.
@@ 4.2 Image Representation
Figure 5: Subjective comparison of various codecs on Kodak at low Bpp.
Table 2: Computational complexity of traditional and learning-based image codecs on DIV2K Dataset at low and high Bpp.
@@
JPEG [61] 0.3197/0.5638 25.2920/28.4299 0.9020/0.9559 608.61/557.35 614.68/545.59
JPEG2000 [55] 0.2394/0.5993 27.2792/30.9294 0.9305/0.9663 3.46/3.40 4.32/3.93
Ballé17 [5] 0.2271/0.4987 27.7168/30.7759 0.9508/0.9775 21.23/16.53 18.83/17.87
Ballé18 [6] 0.2533/0.5415 28.7548/32.2351 0.9584/0.9816 16.53/13.56 15.87/15.20
[…]
@@ 4.3 Image Compression
Computational complexity. Table 2 reports the computational complexity of several image codecs on the DIV2K dataset, with learning-based codecs operating on an NVIDIA V100 GPU and traditional codecs running on an Intel Core(TM) i9-10920X processor at a base frequency of 3.50GHz in single-thread mode. Impressively, the decoding speed of our codec reaches around 2000 FPS, outpacing traditional codecs like JPEG, while also providing enhanced compression performance at lower bitrates. This establishes a significant advancement in the field of neural image codecs.
@@ 4.4 Ablation Study
Quantization strategies. Table 4 investigates the effect of different quantization schemes on image compression. Without the commitment loss $\mathcal{L}_{c}$ (V1), the absence of supervision for the RVQ codebook leads to significant deviations of the codebook vector from the original vector, adversely affecting reconstruction quality. Moreover, eliminating RVQ in favor of 6-bit integer quantization for color parameters (V2) resulted in a 7.02% increase in bitrate consumption when compared with our default solution. This outcome suggests that the color vectors across different Gaussians share similarities, making them more suitable for RVQ. Further exploration into the use of higher bit quantization (V3) reveals a deterioration in compression efficiency.
@@
(V1) w/o $\mathcal{L}_{c}$+w/ RVQ + 6bit -3.145 333.16 -0.0824 337.84
(V2) w/o $\mathcal{L}_{c}$+w/o RVQ + 6bit -0.159 7.02 -0.0030 6.14
(V3) w/o $\mathcal{L}_{c}$+w/o RVQ + 8bit -0.195 11.69 -0.0127 62.77
@@ 5 Conclusion
Acknowledgments. This work was supported by the National Natural Science Fund of China (Project No. 42201461, 62102024, 62331014) and the General Research Fund (Project No. 16209622) from the Hong Kong Research Grants Council.
@@ 0.C.1 Implementation Details
Baselines. For SIREN [54] and WIRE [53], we implement them by using the open-source project [^1] https://github.com/vishwa91/wire from WIRE. For I-NGP [48] and NeuRBF [18], we adopt the project [^2] https://github.com/oppo-us-research/NeuRBF from NeuRBF. As for compression baselines, the implementation of COIN [23] uses their library [^3] https://github.com/EmilienDupont/coin. We evaluate the VAE-based codecs (Ballé17 [5], Ballé18 [6]) using the MSE-optimized models provided by CompressAI [9]. It is worth noting that during the inference of the INR methods, we sample all image coordinates at once to output the corresponding RGB values. This is the maximum inference speed that the INR methods can achieve when the GPU memory resources are sufficient.